The work done by the force F = 4ax - 3ay + 2az N in giving a 1 nC charge a displacement of 10ax + 2ay - 7az m is


1
Expert's answer
2021-07-08T09:58:19-0400

"F=4a_x - 3a_y + 2a_z,\\
l=10a_x + 2a_y - 7a_z,"

"\\A=qFl=q(4a_x - 3a_y + 2a_z)(10a_x + 2a_y - 7a_z)=10^{-9}\cdot(40-6-14)=20\cdot 10^{-9}=20~nJ."